OpenStudy (anonymous):

Can you create an equation for me with an extraneous solution, please? Thanks!

OpenStudy (anonymous):

you mean like one with a square root?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Like sqrt (x - 98) + 2 = 1

OpenStudy (anonymous):

ok lets make it easy pick a number you want to be a solution

OpenStudy (anonymous):

Okay, 4?

OpenStudy (anonymous):

ok so \(\sqrt{4+5}=3\) right? so we will make one side of the equation \[\sqrt{x+5}\] and we know if \(x\) is \(4\) then that side will be \(3\)

OpenStudy (anonymous):

now for the other side \[\sqrt{4+12}=4\] so on the other side we can put \(\sqrt{x+12}-1\) which will also be \(3\) now lets solve \[\sqrt{x+1}=\sqrt{x+12}-1\]

OpenStudy (anonymous):

typo there i meant lets solve \[\sqrt{x+5}=\sqrt{x+12}-1\]

OpenStudy (anonymous):

oh damn i think i messed this one up, will only get one solution

OpenStudy (anonymous):

Oh lol, well would the first step be to square both sides we get x + 5 = x + 12 - 1?

OpenStudy (anonymous):

no, when you square both sides you get \[x+5=x+12-2\sqrt{x+12}+1\]

OpenStudy (anonymous):

lets try this one, might work better if \(x=7\) then \(\sqrt{2x+2}=\sqrt{16}=4\) and \(\sqrt{x+2}=3\) so lets try and solve \[\sqrt{2x+2}=\sqrt{x+2}+1\]

OpenStudy (anonymous):

Wow, I'm sorry, I really suck at math and I'm lost

OpenStudy (anonymous):

ok let me see if i can come up with a really easy example

OpenStudy (anonymous):

Thank you! I really appreciate it.

OpenStudy (zarkon):

\[x-6=\sqrt{x}\]

OpenStudy (anonymous):

much easier